OpenStudy (anonymous):

Can someone please help me with this question? 1. 0031 ^ "x" = 1.66708334 How do I find x? Thank you

OpenStudy (anonymous):

jesus you have to take the ln of both sides to get: ln(1.0031^x)=ln(1.66708334) that simplifies to ln(1.0031)x=ln(1.66708334). Then divide by ln(1.0031) to get: x=ln(1.66708334)/ln(1.0031)

OpenStudy (anonymous):

Thank you so much david.

OpenStudy (anonymous):

What class is that for?

OpenStudy (anonymous):

I'm just reviewing for a Finance exam and it has been years since I did Logarithms. The professor threw a curveball and included logs. I appreciate your help.

OpenStudy (anonymous):

Ok I see, yeah no probs! ;)